# Beyond-Labels: Advancing Open-Vocabulary Segmentation With Vision-Language Models

### v3

@@ Abstract
Self-supervised learning can resolve numerous image or linguistic processing problems when effectively trained. This study investigated simple yet efficient methods for adapting previously learned foundation models for open-vocabulary semantic segmentation tasks. Our research proposed ”BeyondLabels,” a lightweight transformer-based fusion module that uses a handful of image segmentation data to fuse frozen image representations with language concepts. This strategy allows the model to successfully actualize enormous knowledge from pre-trained models without requiring extensive retraining, making the model data-efficient and scalable. Furthermore, we efficiently captured positional information in images using Fourier embeddings, thus improving the generalization across various image sizes, addressing one of the key limitations of previous methods. Extensive ablation tests were performed to investigate the important components of our proposed method; when tested against the common benchmark PASCAL-5i, it demonstrated superior performance despite being trained on frozen image and language characteristics.
@@ I Introduction
Recent research on self-supervised learning has yielded promising results. Methods such as MoCo [6] and DINO [4] have demonstrated that powerful image representations can be learned without utilizing labeled data. Interestingly, even though these models were only trained on unlabeled images, they appear to naturally detect objects in images, allowing for fundamental semantic segmentation [17, 25, 29].
@@
Another technique relies on learning a shared representation for images and text from big image-text pairs available on the internet. CLIP [20] and ALIGN [15], taught using simple contrastive learning techniques, have exhibited impressive capacities to perform new tasks without additional training, such as image classifying in a ”zero-shot” way. As computational power increases, more powerful models are anticipated to emerge, trained on larger datasets, and producing better image categorization results. It has opened a way towards the development of open-vocabulary segmentation, image retrieval, and multimodal reasoning within wide-ranging applications involving vision-language tasks. Larger and more powerful models are envisioned with increases in computational power and larger available datasets for further gains in both accuracy and robustness of image classification, along with other applications involving vision language. The above-mentioned continuous improvement illustrates the increased importance of contrastive learning in closing the gap in visual and textual understanding.
@@
Semantic segmentation is a harder operation involving labeling each image pixel with a certain category. Recent models, such as FCN [24], U-Net [22], DeepLab [5], and DPT [21], have had tremendous success in training deep neural networks on huge datasets created for segmentation. Traditional techniques, however, rely on training classifiers for specific categories, which poses a significant constraint. These models can only predict labels for the categories on which they were trained, therefore, they cannot detect or segment objects from new, previously unseen categories during testing.
@@
In our work, we investigate efficient techniques to connect the powerful pre-trained vision-language models to solve the challenge of open-vocabulary semantic segmentation successfully, ”the segmentation of objects from any category based on their textual names or descriptions.” To capture fine-grained spatial information, we first encode the input image using a frozen, self-supervised image model supplemented with Fourier embeddings, particularly across changing resolutions. Fourier embedding is used to extract high-level spatial/frequency features from image data to help the model learn better. Simultaneously, we encode category names using a frozen language model to ensure a strong representation of textual descriptions. The integrated embeddings from two alternative modalities allow for exact matching of picture regions and textual descriptions, resulting in seamless segmentation across many categories. The key ingredient of our approach is the reliance on Fourier embeddings [27], [13], [18] to encode positional information without resorting to the commonly used pre-trained position embeddings. This enhances the generalization across images and results in smooth and consistent spatial encoding. A transformer-based fusion module receives the combined features. Taking into account data from the other modality, this module uses self-attention to update the features. To create the segmentation mask, we calculate the cosine similarity between each pixel and the text characteristics in each category. The dot product of the L2 normalized image and language embeddings is used to accomplish this.
@@
Among the contributions we provide, this includes a simple, lightweight, portable, transformer-based fusion module, Beyond-Labels, which stitches effectively a series of open-vocabulary pre-trained models. To encode the position, unlike typical usage, common pre-trained position embeddings will not be considered but instead are realized by Fourier embeddings, which bring great improvement in generalization among different images but with smooth, steady spatial encoding guaranteed. We also present ablation studies to verify our approach’s main components. This demonstrates strong zero-shot performance and competitive results on the PASCAL few-shot benchmark.
@@
Figure 1: Our proposed framework consists of a text encoder, a multi-model embedding fusion module, a decoder, and the image encoder utilizing a pre-trained vision-language model with Fourier embeddings. The multimodal embedding fusion module combines the features the frozen encoders have extracted from the text and image data. Before segmenting the image features according to how closely they resemble text features, the decoder upscales them to their original size.
@@ II-A Semantic Segmentation of Unseen Categories
Teaching a network to segment new, previously unknown categories remains difficult, as most existing approaches only function with a predefined set of categories, requiring the test and training categories to match. Zero-shot semantic segmentation attempts to address this issue by employing word embeddings (semantic representations of category names) to recognize new categories without further training data. There are two primary approaches:
  • Generative approaches, such as ZS3Net [3] and CaGNet [12], combine an image segmentation model with a mechanism for producing image features for new categories via word embeddings.
  • Discriminative algorithms, such as SPNet [30] and LSeg [8], map pixels and category names into a shared embedding space, which is then used to forecast the class of each pixel.
II-B Fourier Embedding for Spatial Encoding
Recently, Fourier embeddings have become appealing alternatives to traditional position encodings in those tasks that require strong spatial representation. Unlike learned positional embeddings, which will struggle hard to generalize outside the training distribution, Fourier-based embeddings provide a continuous and smooth encoding of spatial locations. Recently, it has been the case that Fourier embeddings turn out to be useful in improving generalization for implicit neural representations. For example, NeRF [19] proposed using Fourier feature mappings to help the positional encoding for 3D reconstruction tasks. Since then, this idea has been applied to image segmentation, where models benefit from the smoothness and periodicity of Fourier-transformed coordinates. In this context, self-supervised learning for Perceiver IO [14] uses Fourier embeddings to map high-dimensional inputs to a latent representation, with the best performance across various multimodal tasks. Correspondingly, in semantic segmentation, Fourier Network Representations [26] are utilized to encode better spatial dependencies, reducing large labeled datasets. Our approach thus replaces conventional learned embeddings with integrated Fourier embeddings of spatial information for smooth generalization across image domains. This goes in line with previous work showing that continuous and structured position encoding is helpful, both for vision and language models.
II-C Pre-trained Language and Vision Models
@@
In computer vision and natural language processing, self-supervised learning has made substantial progress. For example, methods such as contrastive learning (MoCo [6]), metric learning (BYOL [11]), and self-distillation (DINO [4]) have emerged as effective tools for extracting characteristics for subsequent tasks. Language models, on the other hand, are commonly pre-trained with masked language modeling (BERT [7]) or next-token prediction (GPT [2]).
@@
Large-scale image-language models have recently received increased interest. CLIP [20] is an important example, which trains on massive image-text pairs using contrastive learning and has successfully aligned images and text in a common space. Following that, CLIP’s pre-trained knowledge is applied to various tasks, including object detection, picture captioning, image segmentation, and text-driven image manipulation. Instead of fine-tuning CLIP for each job, we adopt a simpler strategy, integrating a lightweight fusion module with pre-trained self-supervised vision and language models. This methodology is an incredibly effective foundation for the semantic segmentation of new groups.
@@
As foundation models [10, 9, 16] evolve rapidly, numerous studies have investigated how to merge multiple pre-trained models for different tasks. For example, Flamingo [1] proposed an architecture that uses big, pre-trained vision-only and language-only models to perform image-language tasks with minimal data. Socratic Models [31] uses a similar method but integrates pre-trained language models, vision-language models, and audio-language models to do multi-modal tasks like image captioning and video-to-text retrieval without retraining the models. Frozen [28] suggested a method for aligning a pre-trained, frozen language model with an image encoder using only a few examples and continual prompts.
@@ III Methodology
This section begins by defining the open-vocabulary semantic segmentation problem. Next, we describe our suggested architecture, ”Beyond-Labels,” which links the previously trained language and image models to accomplish the task.
@@ III-A Problem Formulation
Where $\mathcal{X}$ is the input picture and $\mathcal{Y}$ is the one-hot encoded segmentation masks for $\mathcal{X}$; $\mathcal{U}$ represents a unit of observed categories in $\mathcal{X}$, and $\mathcal{S}$ represents all training categories in this unit. We aim to train a segmentation model that can separate a test image into significant areas of unexplored categories.
III-B Image Feature Extraction
The process starts with splitting the input image $\mathcal{X}$ as a collection of non-overlapping 2D patches. Each patch is flattened into a vector and put onto an embedding space with dimension ’d’ using a linear projection. Fourier embeddings [27] are used to encode positional information rather than standard pre-trained positional embeddings, allowing for improved generalization to images of varied sizes while also maintaining smooth and consistent spatial encoding. A Fourier embedding is calculated for each patch at point (x,y) on the grid using sinusoidal functions. The Fourier embedding is applied to the patch embedding, resulting in a final position-aware embedding $\mathcal{Z}$ for each patch.
$z_{i}$ $=x_{i}+f\_emb(x,y)$ (2)
Transformer encoder layers process position-aware embeddings, $\mathcal{Z}$ and layer normalization, feedforward networks, multi-head self-attention, and residual connections are used by each layer to analyze the embeddings. CLIP[20], a self-supervised language model trained on a huge collection of texts or images, is used as an image encoder.
$F_{V}$ $=\Phi_{\text{IE}}\left(f\_{\text{emb}}(X_{j})\right)\in\mathbb{R}^{(HW)\times d}$ (3)
Where ”d” is the image embedding dimension, p is the patch size, and H = H/p and W = W/p.
@@ III-C Textual Feature Extraction
Semantic categories ($\mathcal{W}$) are encoded using a text encoder to produce text embeddings.
$F_{W}$ $=\Phi_{\text{TE}}(W)\in\mathbb{R}^{|W|\times d}$ (4)
@@
Where ”d” is the dimension of the word embeddings and ”W” is the number of input categories. We employ prompt templates such as ”a picture of a category within the context” to provide semantic category context. The output embeddings generated by the text encoder are then averaged. The ”Result Section” offers a collection of prompt templates. We use CLIP [20], a self-supervised language model, as a text encoder.
@@
Figure 2: Overview of the decoder section. Multimodal embedding fusion combines image and text encoder features that are processed through a decoder to generate upsampled image features. These were then compared with the textual features using cosine similarity, creating a segmentation mask with labels refined by thresholding and a temperature-scaled sigmoid.
III-D Multi_Model Embedding Fusion
@@
To align the image and textual information, we use MLPs ($\theta$) to align their channel dimensions ($F_{V}$ and $F_{W}$). To capture the interactions between the textual and image data, we subsequently pass the combined features through a multimodal embedding fusion module:
$[F_{V}^{0},F_{W}^{0}]=\Phi_{\text{Fuse}}(\theta[F_{V},F_{W}])$ (5)
@@
The fusion module uses multiple transformer encoder layers to capture image-text relationships. The fused image features ($F_{V}^{0}$) and text features ($F_{W}^{0}$) resemble the originals ($F_{V}$) and ($F_{W}$) in shape, but they are improved and enhanced by constantly focusing on one another’s modality. This enrichment comes because the fusion module iteratively puts more focus on cross-modal interactions, which allow the features of the image and text modalities to enrich each other’s representations. Thus, the fused features become better aligned and may be capable of providing a more holistic representation of the underlying data, therefore boosting performance in further tasks involving either images or text. Such an attention-driven fusion process significantly enhances the model’s capability of capturing fine-grained correlations between the two modalities, which is crucial for achieving high-quality results on tasks like open-vocabulary semantic segmentation. Our approach exploits the complementary strengths of both image and text data to allow more accurate and context-aware embeddings
III-E The Hierarchical Decoder
@@
The multimodal embedding fusion combines the image and text encoders’ features. These combined features are passed to the decoder for further processing. The decoder gives us the image feature map, which is progressively upsampled to match the resolution of the input image. The upsampled image features are compared with the textual features using cosine similarity to generate the logits ŷ. This comparison helps determine how closely each area in the image aligns with the provided text categories. The final result is a segmentation mask, where different regions of the image are assigned labels. Using class-wise thresholding and a sigmoid with a temperature ($\mathcal{\tau}$), final predictions can be made using binary segmentation for each category.
@@ IV Experimental Results
Model 5_0 5_1 5_2 5_3 mIoU
SPNet [30] 23.8 17.0 14.1 18.3 18.3
ZS3Net [3] 40.8 39.4 39.3 33.6 38.3
OSLSM [23] 33.6 55.3 40.9 33.5 40.8
Ours 44.6 49.4 40.0 32.1 41.5
TABLE I: Performances of different methods on the evaluation dataset of PASCAL-5i.
Model F_Emb Fusion Image_Decoder 50 51 52 53 mIoU
B_L_0 X X 32.6 34.1 25.2 21.0 25.7
B_L_1 X 38.4 44.6 35.4 29.3 36.9
B_L_2 44.6 49.4 40.0 32.1 41.5
TABLE II: Ablation study on PASCAL-5i

### v5

@@ Abstract
Open-vocabulary semantic segmentation attempts to classify and outline objects in an image using arbitrary text labels, including those unseen during training. Self-supervised learning resolves numerous visual and linguistic processing problems when effectively trained. This study investigates simple yet efficient methods for adapting previously learned foundation models for open-vocabulary semantic segmentation tasks. Our research proposes ”Beyond-Labels,” a lightweight transformer-based fusion module that uses a handful of image segmentation data to fuse frozen visual representations with language concepts. This strategy allows the model to successfully actualize enormous knowledge from pre-trained models without requiring extensive retraining, making the model data-efficient and scalable. Furthermore, we efficiently capture positional information in images using Fourier embeddings, thus improving the generalization and resulting in smooth and consistent spatial encoding. We perform thorough ablation studies to investigate the major components of our proposed method in comparison to the standard benchmark PASCAL-5i, the method performs better despite being trained on frozen vision and language characteristics.
@@ I Introduction
Recent research on segmentation has yielded promising results. Methods such as MoCo [5] and DINO [4] have demonstrated that powerful visual representations can be learned without utilizing labeled data. Interestingly, even though these models were only trained on unlabeled images, they appear to naturally detect objects in images, allowing for fundamental semantic segmentation [15, 23].
@@
A recent trend in segmentation is to combine vision and text domains. CLIP [17] and ALIGN [14], taught using simple contrastive learning techniques, have exhibited impressive capacities to perform new tasks without additional training, such as image classifying in a ”zero-shot” way. As computational power increases, more powerful models are anticipated to emerge, trained on larger datasets and producing better image categorization results. It has opened a way towards the development of open-vocabulary segmentation, image retrieval, and multimodal reasoning within wide-ranging applications involving vision-language tasks. Larger and more powerful models are envisioned with increases in computational power and larger available datasets for further gains in both accuracy and robustness of image classification, along with other applications involving vision language. The above-mentioned continuous improvement illustrates the increased importance of contrastive learning in closing the gap in visual and textual understanding.
@@
Semantic segmentation is a harder operation involving labeling each image pixel with a certain category. Recent models, such as FCN [22], U-Net [20], and DPT [19], have had tremendous success in training deep neural networks on huge datasets created for segmentation. Traditional techniques, however, rely on training classifiers for specific categories, which poses a significant constraint. These models can only predict labels for the categories on which they were taught; therefore, they cannot detect or segment objects from new, previously unseen categories during testing.
@@
In our work, we investigate efficient techniques to connect the powerful pre-trained vision-language models to solve the challenge of open-vocabulary semantic segmentation successfully, ”the segmentation of objects from any category based on their textual names or descriptions.” To capture fine-grained spatial information, we first encode the input image using a frozen, self-supervised vision model supplemented with positional encoding. Fourier embedding is used to extract high-level spatial/frequency features from image data to help the model learn better. Simultaneously, we encode category names using a frozen language model to ensure a strong representation of textual descriptions. The integrated embeddings from two alternative modalities allow for exact matching of picture regions and textual descriptions, resulting in seamless segmentation across many categories. The key ingredient of our approach is the reliance on Fourier embeddings [26] to encode positional information without resorting to the commonly used pre-trained position embeddings. This enhances image generalization and results in smooth and consistent spatial encoding. A transformer-based [28] fusion module receives the combined features. Taking into account data from the other modality, this module uses self-attention to update the features. To create the segmentation mask, we calculate the cosine similarity between each pixel and the text characteristics in each category. The dot product of the L2-normalized vision and language embeddings is used to accomplish this.
@@
Among the contributions we provide, this includes a simple, lightweight, portable, transformer-based fusion module, Beyond-Labels, which stitches together effectively a series of open-vocabulary pre-trained models. To encode the position, unlike typical usage, common pre-trained position embeddings were not considered but instead are realized by Fourier embeddings, which bring great improvement in generalization among different images, but with smooth, steady spatial encoding guaranteed. This demonstrates strong zero-shot performance and competitive results on the PASCAL few-shot benchmark.
@@
Figure 1: Our proposed framework consists of a text encoder, a multimodal embedding fusion module, a decoder, and the vision encoder utilizing a pre-trained vision-language model with Fourier embeddings. The multimodal embedding fusion module combines the features that the frozen encoders have extracted from the text and image data. Before segmenting the visual features according to how closely they resemble text features, the decoder upscales them to their original size.
@@ II Related Work
We start by discussing semantic segmentation of novel classes, where word embeddings are used by zero-shot learning approaches to label unseen objects without training on extra data. Then, we talk about pre-trained language and vision models, wherein we present large-scale self-supervised models like CLIP that align images and text in the same vector space. Finally, we move on to Fourier embeddings for spatial encoding, which improves generalization by giving smooth, continuous positional representations for applications like image segmentation.
@@ II-A Semantic Segmentation of Unseen Categories
Teaching a network to segment new, previously unknown categories remains difficult, as most existing approaches only function with a predefined set of categories, requiring the test and training categories to match. Zero-shot semantic segmentation attempts to address this issue by employing word embeddings (semantic representations of category names) to recognize new categories without further training data. There are two primary approaches: Generative approaches, such as ZS3Net [3] and CaGNet [11], which combine an image segmentation model with a mechanism for producing visual features for new categories via word embeddings, Discriminative algorithms, such as SPNet [30] map pixels and category names into a shared embedding space, which is then used to forecast the class of each pixel.
II-B Pre-trained Language and Vision Models
@@
Self-supervised learning has made substantial progress in computer vision and natural language processing. For example, methods such as contrastive learning (MoCo [5]), metric learning (BYOL [10]), and self-distillation (DINO [4]) have emerged as effective tools for extracting characteristics for subsequent tasks. Language models, on the other hand, are commonly pre-trained with masked language modeling (BERT [6]) or next-token prediction (GPT [2]).
@@
Large-scale vision-language models have recently received increased interest. CLIP [17] is an important example, which trains on massive image-text pairs using contrastive learning and has successfully aligned images and text in a common space. Following that, CLIP’s pre-trained knowledge is applied to various tasks, including object detection, picture captioning, image segmentation, and text-driven image manipulation. Instead of fine-tuning a pre-trained mode for each job, we adopt a simpler strategy, integrating a lightweight fusion module with pre-trained self-supervised vision and language models. This methodology is an incredibly effective foundation for the semantic segmentation of new groups.
@@
As foundation models [8, 7] evolve rapidly, numerous studies have investigated how to merge multiple pre-trained models for different tasks. For example, Flamingo [1] proposed an architecture that uses big, pre-trained vision-only and language-only models to perform image-language tasks with minimal data. Socratic Models [31] uses a similar method but integrates pre-trained language models, vision-language models, and audio-language models to do multi-modal tasks like image captioning and video-to-text retrieval without retraining the models. Frozen [27] suggested a method for aligning a pre-trained, frozen language model with a visual encoder using only a few examples and continual prompts.
II-C Fourier Embedding for Positional Encoding
Recently, Fourier embeddings have become appealing alternatives to traditional position encodings in those tasks that require a strong spatial representation. Unlike learned positional embeddings, which will struggle hard to generalize outside the training distribution, Fourier-based embeddings provide a continuous and smooth encoding of spatial locations. Fourier embeddings turn out to be useful in improving generalization for implicit neural representations. For example, NeRF [16] proposed Fourier feature mappings to help the positional encoding for 3D reconstruction tasks. Self-supervised learning for Perceiver IO [13] uses Fourier embeddings to map high-dimensional inputs to a latent representation, with the best performance across various multimodal tasks. In semantic segmentation, Fourier network representations [25] are utilized to encode better spatial dependencies, reducing large labeled datasets. Our approach thus replaces conventional learned embeddings with integrated Fourier embeddings of spatial information for smooth generalization across image domains. This goes in line with previous work showing that continuous and structured position encoding is helpful, both for vision and language models.
@@ III Methodology
This section begins by defining the open-vocabulary semantic segmentation problem. Next, we describe our suggested architecture, ”Beyond-Labels,” given in Fig 1, which links the previously trained language and vision models to accomplish the task.
@@ III-A Problem Formulation
Where $\mathcal{X}$ is the input picture, and $\mathcal{Y}$ is the one-hot encoded segmentation mask for $\mathcal{X}$; $\mathcal{U}$ represents a unit of observed categories in $\mathcal{X}$, and $\mathcal{S}$ represents all training categories in this unit. We aim to train a segmentation model that separates a test image into significant areas of unexplored categories.
III-B Visual Feature Extraction
The process starts with splitting the input image $\mathcal{X}$ as a collection of non-overlapping 2D patches. Each patch is flattened into a vector and put onto an embedding space with dimension ’d’ using a linear projection. Fourier embeddings [26] are used to encode positional information rather than standard pre-trained positional embeddings, allowing for improved generalization to images while also maintaining smooth and consistent spatial encoding. Using sinusoidal functions, a f-emb is calculated for each patch at point (x, y) on the grid. The Fourier embedding is applied to the patch embedding, resulting in a final position-aware embedding for each patch.
$X_{i}$ $=x_{i}+f\_emb(x,y)$ (2)
Here $x_{i}$ represents the flattened vector of the image patch. CLIP[17], a self-supervised language model, is used as a visual encoder.
[…]
Here ’d’ is the visual embedding dimension, $\mathcal{H}=\frac{\mathcal{H}}{p}$ and $\mathcal{W}=\frac{\mathcal{W}}{p}$ where $p$ is the patch size.
@@ III-C Textual Feature Extraction
Semantic categories ($W$) are encoded using a text encoder to produce text embeddings.
[…]
@@
Where ’d’ is the dimension of the word embeddings and ’W’ is the number of input categories. We employ prompt templates such as ”a picture of a category within the context” to provide semantic category context. The output embeddings generated by the text encoder are then averaged. The ”Appendix Section” offers a collection of prompt templates. We use CLIP [17], a self-supervised language model, as a text encoder.
@@
Figure 2: Overview of the decoder section. Multimodal embedding fusion combines visual and text encoder features that are processed through a decoder to generate upsampled visual features. These are then compared with the textual features using cosine similarity, creating a segmentation mask with labels refined by thresholding and a temperature-scaled sigmoid.
III-D MultiModal Embedding Fusion
@@
To align the visual and textual information, we use MLPs ($\theta$) to align their channel dimensions ($F_{i}$ and $F_{W}$). To capture the interactions between the textual and visual data, we subsequently pass the combined features through a multimodal embedding fusion module:
$[F_{i\_upd},F_{w\_upd}]=\textit{Fuse}(\theta[F_{i},F_{w}])$ (5)
@@
The fusion module uses multiple transformer encoder layers to capture visual-text relationships. The fused visual $F_{i\_upd}$ and text features $F_{w\_upd}$ resemble the originals ($F_{i}$) and ($F_{w}$) in shape, but they are improved and enhanced by constantly focusing on one another’s modality. This enrichment comes because the fusion module iteratively puts more focus on cross-modal interactions, which allow the features of the visual and text modalities to enrich each other’s representations. Thus, the fused features become better aligned and may provide a more holistic representation of the underlying data, boosting performance in further tasks involving visuals or text. Such an attention-driven fusion process significantly enhances the model’s capability of capturing fine-grained correlations between the two modalities, which is crucial for achieving high-quality results on tasks like open-vocabulary semantic segmentation. Our approach exploits the complementary strengths of both visual and text data to allow more accurate and context-aware embeddings
III-E The Hierarchical Visual Decoder
@@
The multimodal embedding fusion combines the visual and text encoders’ features. These combined features are passed to the Hierarchical Visual Decoder for further processing. The decoder gives us the visual feature map, which is progressively upsampled to match the resolution of the input image. The upsampled visual features are compared with the textual features using cosine similarity to generate the logits ŷ. This comparison helps determine how closely each area in the image aligns with the provided text categories. The final result is a segmentation mask, where different regions of the image are assigned labels. Using class-wise thresholding and a sigmoid with a temperature ($\mathcal{\tau}$), final predictions are made using binary segmentation for each category.
@@ IV Experimental Results
Our experiment section covers the PASCAL-5i dataset for few-shot segmentation evaluation, the evaluation Metric, the performance comparison across models, and the contribution of the different components of the proposed model.
TABLE I: Performance comparison of different methods on the evaluation dataset of PASCAL-5i.
Model Method 50 51 52 53 mIoU
LogReg [21] 1-shot 26.9 42.9 37.1 18.4 31.4
Siamese [21] 1-shot 28.1 39.9 31.8 25.8 31.4
Finetuning [21] 1-shot 24.9 38.8 36.5 30.1 32.6
1-NN [21] 1-shot 25.3 44.9 41.7 18.4 32.6
OSLSM [21] 1-shot 33.6 55.3 40.9 33.5 40.8
CO-FCN [18] 1-shot 36.7 50.6 44.9 32.4 41.1
AMP-2 [24] 1-shot 41.9 50.2 46.7 34.7 43.4
SG-One [32] 1-shot 40.2 58.4 48.4 38.4 46.3
PANet [29] 1-shot 44.0 57.5 50.8 44.0 49.1
SPNet [30] zero-shot 23.8 17.0 14.1 18.3 18.3
ZS3Net [3] zero-shot 40.8 39.4 39.3 33.6 38.3
Ours zero-shot 46.6 56.2 51.6 43.2 49.4
TABLE II: Table representing object categories for different values of PASCAL-5i.
i = 0 i = 1 i = 2 i = 3
aeroplane, bicycle, bird, boat, bottle bus, car, cat, chair, cow diningtable, dog, horse, motor-bike, person potted plant, sheep, sofa, train, tv/monitor
TABLE III: Ablation study on PASCAL-5i. The bold indicates the best results.
Model F_Emb Fusion Visual_Decoder 50 51 52 53 mIoU
B_L_0 X X 36.6 43.1 35.2 27.0 34.0
B_L_1 X 40.4 46.7 38.4 32.3 39.5
B_L_2 46.6 56.2 51.6 43.2 49.4